OpenStudy (anonymous):

How do you solve the problem Pleaseeeeee help The number of drive-in movie screens has declined steadily in the united states. Pennslyvania and Ohio had the most remaining drive-in movie screens in 2007, Pennsylvania had 2 more screens than Ohio, and there were 68 screens total in the two states. How many drive- in movie screens remained In each state?

OpenStudy (anonymous):

Let's use abbreviations to make this easier to write and read: Ohio (OH) and Pennsylvania (PA) PA has 2 more than OH, so PA = OH + 2 The total (68) is PA and OH together, so PA + OH = 68 Now you need to use both pieces of information to find the answer in terms of OH: PA + OH = (OH + 2) + OH = 2 OH + 2 = 68 Can you take it from there?

OpenStudy (anonymous):

Sorry I'm really stupid but I don't really understand it

OpenStudy (anonymous):

Do you understand how I got what I did so far, or do you need help with going on from there?